OpenStudy (anonymous):

Medal for answer! Find all real solutions of the equation. (Enter your answers as a comma-separated list. If there is no real solution, enter NO REAL SOLUTION.) 8t^10 − 3t^7 = 0

OpenStudy (anonymous):

t^7(8t^3-3)=0 ---->t=0 or 8t^3=3 t^3=3/8 ---->t=cube root of 3/8